# Information Access of the Oppressed: A Problem-Posing Framework for Envisioning Emancipatory Information Access Platforms

### v1

@@ 6. Conclusion
Information access is too critical to our collective sensemaking and sociopolitical discourse to be left in the hands of powerful corporations and economic elites. In a globalized world, our IA platforms must represent the interconnectedness of different communities and cultures worldwide, and not privilege the perspectives of a handful of powerful nations nor continue to reproduce the epistemic injustices of our past. These injustices are the poison fruits of the structures of oppression that have their roots in our history and continue their domination (Collins, 1990) into the present, including colonialism, cisheteropatriarchy, classism, casteism and ableism among others; and so it is the explicit goal of emancipatory IA to aid the dismantling of these structures.
@@
Our approach to emancipatory IA is predicated on a rejection of technosolutionism in that it is not the design of the technology that confronts oppression; but rather it is the creation of spaces for community participation, collective deliberation, and democratic negotiation within the platform, that is so often missing in our modern technology design, that we believe holds the key to realizing emancipatory access to information. The role of design is to structurally encode these spaces within the IA platform, which is necessary but in itself not sufficient for the platform to be emancipatory. Emancipation after all is a political project and so it is the social arrangements within these spaces that are key to determining whether the platform will be emancipatory or oppressive. Emancipatory IA does not brush this political reality under the rug but rather foregrounds it. To bend the arc of online information access towards emancipation we need to experiment with different structures of democratic participation, governance, and decision-making. But beyond structures and processes, we also need to ground these discourses and negotiations in critical scholarship, theories, and practices that distinguishes structural understanding of systems of oppression from reactionary narratives. For this, we must lean on relevant disciplines such as critical theory, information and library sciences, science and technology studies, and others; while being clear-eyed about the growing phenomenon of reactionary co-option and weaponization of social justice language to prop up false victimhood—e.g., (Strickland, 2016; Center, ; Savera, 2024). Our “neutrality” must manifest through equal treatment of all groups with respect to humanization and freedom from oppression; not drawing of false equivalences between oppressors and oppressed.
@@
Our problem-posing framework, as the name explicitly suggests, does not point to miracle-cures for the ailing growth of global authoritarianism and oppression. Instead, it responds to the call for building an emancipatory movement within the IA community (Mitra, 2025a) by identifying some of the key challenges that the movement should focus on. This problem-posing framework is intended to be a concrete starting point for envisioning radically new alternatives. Envisioning in this context must not be idle conjectures but praxis; we must use these questions to reflect and act on the problems collectively and in the process co-construct the social arrangements of individuals and institutions committed to the causes of universal humanization and emancipation.

### v2

@@ 6. Conclusion
IA is too critical to our collective sensemaking and sociopolitical discourse to be left in the hands of powerful corporations and economic elites. In a globalized world, our IA platforms should represent the interconnectedness of different communities and cultures worldwide, and not privilege the perspectives of a few powerful nations nor continue to reproduce the epistemic injustices of our past. These injustices are the poison fruits of systemic oppression that have their roots in our history and continue their domination (Collins, 1990) into the present—e.g., colonialism, cisheteropatriarchy, classism, casteism, and ableism; and so it is the explicit goal of emancipatory IA to dismantle these structures.
@@
Our approach to emancipatory IA is predicated on the rejection of technosolutionism. It is not the design of the technology that confronts oppression, but rather it is the creation of spaces for community participation, collective deliberation, and democratic negotiation within the platform—that is so often missing in our modern technology design—that we believe holds the key to realizing emancipatory access to information. The role of design is to structurally encode these spaces within the IA platform, which is necessary but in itself not sufficient for the platform to be emancipatory. Emancipation after all is a political project and so it is the sociopolitical arrangements within these spaces that are key to determining whether the platform will be emancipatory or oppressive. Emancipatory IA does not brush this political reality under the rug but rather foregrounds it. To bend the arc of IA towards emancipation we need to experiment with different structures of democratic participation, governance, and decision-making. But beyond structures and processes, we also need to ground these discourses and negotiations in critical scholarship, theories, and practices that distinguish structural understanding of systems of oppression from reactionary narratives. For this, we should lean on relevant disciplines such as critical theory, library and information sciences, science and technology studies, and others; while being clear-eyed about the growing phenomenon of reactionary co-option and weaponization of social justice language to prop up false victimhood—e.g., (Strickland, 2016; Center, ; Savera, 2024). Our “neutrality” should manifest through equal treatment of all groups with respect to humanization and emancipation; not drawing of false equivalences between oppressors and oppressed.
@@
Our problem-posing framework, as the name explicitly suggests, does not point to miracle-cures for the ailing growth of global authoritarianism and oppression. Instead, it responds to the call for building an emancipatory movement within the IA community (Mitra, 2025a) by identifying some of the key challenges that the movement should focus on. This problem-posing framework is intended to be a concrete starting point for envisioning radically new alternatives. Envisioning in this context should not be idle conjectures but praxis; we should use these questions to reflect and act on the problems collectively and in the process co-construct the social arrangements of individuals and institutions committed to the causes of universal humanization and emancipation.